# REArtGS++: Generalizable Articulation Reconstruction with Temporal Geometry Constraint via Planar Gaussian Splatting

### v2

@@ Abstract
Articulated objects are pervasive in daily environments, such as drawers and refrigerators. Towards their part-level surface reconstruction and joint parameter estimation, REArtGS [wu2025reartgs] introduces a category-agnostic approach using multi-view RGB images at two different states. However, we observe that REArtGS still struggles with screw-joint or multi-part objects and lacks geometric constraints for unseen states. In this paper, we propose REArtGS++, a novel method towards generalizable articulated object reconstruction with temporal geometry constraint and planar Gaussian splatting. We first model a decoupled screw motion for each joint without type prior, and jointly optimize part-aware Gaussians with joint parameters through part motion blending. To introduce time-continuous geometric constraint for articulated modeling, we encourage Gaussians to be planar and propose a temporally consistent regularization between planar normal and depth through Taylor first-order expansion. Extensive experiments on both synthetic and real-world articulated objects demonstrate our superiority in generalizable part-level surface reconstruction and joint parameter estimation, compared to existing approaches. Project Site: https://sites.google.com/view/reartgs2/home.
@@ 1 Introduction
Articulated objects are ubiquitous in daily life, such as drawers and laptops, and their surface reconstruction holds significant value across various fields, including robotics and embodied intelligence [wang2025adamanip, Yu2025ArtGS3DGS, yu2024gamma]. Unlike static objects, articulated objects consist of multiple parts with diverse kinematic structures, requiring additional part-level mesh reconstruction and joint parameter estimation. Under this circumstance, ASDF [ASDF] and DITTO [jiang2022ditto] train generative models of articulated objects using expensive 3D supervisions. However, these methods exhibit limited generalization to unseen categories due to the variations in shape and motion of articulated objects. To address this limitation, PARIS [jiayi2023paris] achieves category-agnostic articulated modeling, i.e., part-level reconstruction and motion analysis, using neural implicit radiance fields, only with multi-view RGB images from two different states. Nevertheless, PARIS faces challenges in multi-part perception and smooth dynamic surface reconstruction, suffering from the entangled implicit representation.
@@
Recently, ArtGS [liu2025building] and REArtGS [wu2025reartgs] achieve efficient and realistic articulated object reconstruction and joint parameter estimation from two-state RGB-D and RGB images respectively, leveraging the explicit representation of 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians]. However, we observe that these methods still suffer from the following limitations: (1) They rely on joint type priors, which require additional pipelines and pre-defined thresholds to obtain, and the joint types are assumed to be either rotation or translation. It may lead to significant errors when dealing with screw-joint and multi-part objects; (2) It is difficult to impose temporal geometric constraints from only two-state supervisions. Since the joint parameters and Gaussian primitives are jointly optimized, lacking geometric constraint for unseen states futher affects joint parameter estimation. Although REArtGS enhances geometric constraints by incorporating signed distance fields (SDF), it fails to perform temporal regularization due to the static mapping of SDF.
@@
To introduce dynamic geometric constraint for improved optimization of articulated modeling, we first compress the 3D Gaussians to flat planes to obtain accurate normals and unbiased depth maps following PGSR [pgsr]. Then we propose a temporal geometric constraint, extending the consistent regularization of normal and depth from discrete states to entire time interval through Taylor first-order expansion. Moreover, we propose local consistent voting to address the ambiguous segmentation of overlapping regions. Experimental results of both synthetic and real-world data demonstrate our superiority over existing state-of-the-art (SOTA) methods.
@@ 2.1 Dynamic Reconstruction and Segmentation via 3DGS
3D Gaussian Spaltting [kerbl3Dgaussians] and its subsequent works [pgsr, Huang2DGS2024, yu2024gsdf, scaffoldgs] achieve realistic and efficient scene reconstruction via explicit 3D Gaussian primitives. Recent studies [yang2023gs4d] have increasingly explored the potential of 3D Gaussian Splatting (3DGS) for dynamic scene reconstruction. Specifically, Deformable-3DGS [kerbl3Dgaussians] and 4DGS [4dgs] leverage deformation networks alongside 3D Gaussians, obtaining time-independent Gaussians for dynamic scenes. SC-GS [huang2023sc] and SP-GS [spgs] employ superpoints for dynamic Gaussian modeling and control. Although yielding promising results, each Gaussian learns independent motion in these methods, leading to inconsistency within a rigid part when dealing with articulated objects. Moreover, they require time-continuous supervisions and face challenges in inferring motion using image collections of only two states.
Several works attempt to perform instance segmentation with 3D Gaussians [zhang2025cob, qin2023langsplat, shi2024language, votesplat, shorinwa2024fast]. These methods achieve 3D Gaussian segmentation by introducing 2D masks obtained from foundation models such as SAM [kirillov2023segany] or DINO [zhang2022dino]. However, they rely on the prior knowledge provided by external models and typically struggle with part segmentation.
@@ 2.2 Articulated Object Reconstruction and Motion Analysis
Previous methods [ASDF, song2024reacto] mainly focus on dynamic reconstruction of articulated objects. To achieve both surface reconstruction and motion analysis for articulated objects, PARIS [jiayi2023paris, weng2024neural] employ composite neural radiance fields to jointly optimize geometry and motion through multi-view RGB images of two different states. Following this setting, several methods [guo2025articulatedgs, shen2025gaussianart, liu2025building, wu2025reartgs] introduce 3DGS to articulated object reconstruction and joint parameter estimation. However, these methods require additional pipelines to obtain joint type priors, which are limited to rotation or translation, and affect the performance when dealing with screw-joint or multi-part objects.
@@ 3 Preliminary: Planar Gaussian Spaltting
3DGS [kerbl3Dgaussians] uses a set of Gaussians to explicitly represent the scene instead of neural implicit representation. The rendering color can be derived from the $\alpha$-blending as follows:
@@
where $\mathcal{L}_{1}$ is a pixel-level L1 norm and $\mathcal{L}_{\text{D-SSIM}}$ is the D-SSIM loss [kerbl3Dgaussians].
Recently, PGSR [pgsr] obtains accurate normals and unbiased depth estimates by compressing 3D Gaussians to 2D planar representations, enhancing geometric learning. Concretely, the normal maps $\mathbf{N}$ of planar Gaussians can be naturally rendered via their shortest axes and view directions, and the unbiased depth can be estimated as:
@@ 4.1 Part-Aware Planar Gaussian Representation
Subsequently, we introduce a segmentation mask $\mathbf{M}=\left\{m_{1},m_{2},...,m_{k}\right\}$ for each planar Gaussian, representing the probability belonging to each part. Intuitively, the probability that a Gaussian primitive $\mathcal{G}_{i}$ belongs to a part is inversely proportional to its distance from the part’s center. Inspired by REACTO [song2024reacto], we first denote part centers as learnable parameters $(\mathbf{O},\mathbf{V},\mathbf{\Lambda})$, where $\mathbf{O}\in\mathbb{R}^{k\times 3}$, $\mathbf{V}\in\mathbb{R}^{k\times 3\times 3}$, $\mathbf{\Lambda}\in\mathbb{R}^{k\times 3}$ are part centers, center orientations and scale vectors respectively. For Gaussian $\mathcal{G}_{i}$, the relative distances $\mathbf{L}_{i}$ between Gaussian center $\mu_{i}$ and part centers are defined as: $\mathbf{L}_{i}=\mathbf{V}(\mu_{i}-\mathbf{O})\mathbf{\Lambda}$. Then the Mahalanobis distances $\mathbf{\gamma}$ can be derived as: $\mathbf{\gamma}_{i}=\mathbf{L}_{i}^{T}\mathbf{L}_{i}$. We calculate part segmentation mask $\mathbf{M}_{i}$ of $\mathcal{G}_{i}$ through the Mahalanobis distances and a residual term learned by an multi-layer perceptron (MLP), which is illustrated in the supplementary materials.
@@ 4.2 Part-level Blending Motion
Table 1: Quantitative results on PARIS dataset. We implement all methods without depth supervision for fair comparison. Axis Pos results are measured by cm. ”-” indicates the object containing only prismatic joints. We highlight best and second best results.
@@
Axis Ang PARIS [jiayi2023paris] 15.72 2.02 0.03 7.43 16.62 7.17 0.71 15.40 38.28 0.03 10.34 2.90 20.16 11.53
DTA [weng2024neural] 0.03 0.08 0.07 0.21 0.10 0.08 0.11 0.38 0.18 0.09 0.13 2.26 10.57 6.42
ArtGS [liu2025building] 0.01 0.03 <<0.01 0.02 <<0.01 0.01 0.02 0.04 0.02 0.01 0.02 2.03 4.56 3.30
REArtGS [wu2025reartgs] 0.01 0.02 0.02 0.07 0.05 0.06 0.04 0.08 0.04 0.04 0.04 2.19 87.51 44.85
@@
Axis Pos PARIS [jiayi2023paris] 10.30 9.18 0.76 5.15 15.31 35.07 13.26 24.14 - - 14.15 50.28 - 50.28
DTA [weng2024neural] 1.47 1.16 2.32 1.18 1.95 1.67 1.46 4.87 - - 2.01 59.13 - 59.13
ArtGS [liu2025building] 0.06 0.43 0.94 1.35 0.18 0.38 0.15 1.06 - - 0.51 42.83 - 42.83
REArtGS [wu2025reartgs] 6.24 0.15 1.31 3.64 0.06 0.72 1.56 0.04 - - 1.72 45.72 - 45.72
@@
Part Motion PARIS [jiayi2023paris] 45.80 68.21 0.03 9.70 71.42 128.73 34.57 43.58 0.45 0.30 40.28 2.23 0.57 1.40
DTA [weng2024neural] 0.10 0.12 0.11 0.12 0.34 0.08 0.15 0.24 <<0.01 <<0.01 0.13 2.56 0.18 1.37
ArtGS [liu2025building] 0.03 0.03 0.03 0.01 0.01 0.03 <<0.01 0.02 <<0.01 0.01 0.02 2.27 0.15 1.21
REArtGS [wu2025reartgs] 0.31 0.23 0.06 0.08 0.06 0.15 0.05 0.04 <<0.01 0.08 0.11 38.38 0.52 19.45
@@
CD-s PARIS [jiayi2023paris] 8.51 6.30 0.25 4.57 12.71 2.85 1.95 8.37 10.64 8.62 6.48 10.74 27.61 19.18
DTA [weng2024neural] 0.27 2.48 0.64 10.06 1.64 3.08 2.64 20.73 0.46 10.73 5.74 4.50 30.98 17.74
ArtGS [liu2025building] 0.16 1.89 0.80 8.58 0.50 3.55 2.51 23.63 0.51 11.82 5.40 3.39 18.19 10.79
REArtGS [wu2025reartgs] 0.31 1.53 0.48 6.59 0.25 1.47 0.82 16.04 0.69 7.47 3.57 7.67 22.53 15.10
@@
CD-m PARIS [jiayi2023paris] 4.05 1.26 0.15 3.70 10.06 1.18 3.02 1.52 5.13 20.67 5.07 74.20 314.37 194.29
DTA [weng2024neural] 0.46 2.27 0.96 1.36 10.25 1.61 1.37 4.45 2.07 5.34 3.01 3.63 81.36 42.50
ArtGS [liu2025building] 0.52 1.40 0.98 2.52 0.70 0.86 1.15 4.67 1.23 6.56 2.06 3.23 33.65 18.44
REArtGS [wu2025reartgs] 0.30 0.66 0.71 1.25 0.25 0.21 0.40 0.42 1.97 12.39 1.85 2.87 175.78 89.33
@@
CD-w PARIS [jiayi2023paris] 8.27 2.83 3.13 11.08 11.68 2.80 12.46 19.67 0.81 8.37 8.11 8.82 24.73 16.78
DTA [weng2024neural] 0.58 2.36 0.75 9.35 0.65 2.95 1.43 18.68 0.51 8.34 4.56 4.47 23.48 13.98
ArtGS [liu2025building] 0.44 2.01 0.96 8.32 0.72 2.71 1.57 21.69 0.67 8.69 4.78 3.26 12.26 15.52
REArtGS [wu2025reartgs] 0.70 1.64 0.65 10.89 0.31 3.26 1.54 17.82 0.40 7.71 4.49 2.93 11.48 7.21
@@ 4.5 Initialization and Optimization
Inspired by ArtGS, we first perform a fast optimization for two sets of Gaussians at both start and end states using vanilla 3DGS pipeline [kerbl3Dgaussians], then identify the dynamic Gaussians for each state by criterion: $\Delta\textbf{x}_{i}>\tau\textbf{x}$, where $\Delta\textbf{x}_{i}$ is the Chamfer Distance (CD) from Gaussian $\mathcal{G}_{i}$ at one state to all Gaussians of another state. x is the maximum CD and $\tau$ is set to 0.02. We initialize static and dynamic Gaussians of canonical state as the mean of the two states, using Hungarian matching.
@@ 4.6 Dynamic Mesh Extraction
Table 2: Quantitative results on ArtGS-Multi dataset with additional 2 objects exhibiting screw joints. Due to the large number of parts, we report the average metric for all movable parts. We implement all methods without depth supervision for fair comparison. Part motion error for screw objects is decomposed by translation||rotation. Axis Pos results are measured by cm. ‘-’ indicates the object contains only prismatic joints. We highlight best and second best results.
@@
Axis Ang ArtGS [liu2025building] 1.23 83.69 0.01 34.83 5.62 25.08 42.28 87.57 64.93
REArtGS [wu2025reartgs] 85.45 42.18 64.34 30.69 24.50 49.43 46.45 87.80 67.13
@@
Axis Pos ArtGS [liu2025building] - 70.93 0.36 235.19 8.54 96.49 50.64 38.75 44.70
REArtGS [wu2025reartgs] - 155.80 500.51 31.44 233.48 230.31 47.18 20.41 33.80
@@
Part Motion ArtGS [liu2025building] 0.08 22.50 0.01 18.75 5.73 9.41 0.06||65.87 0.02||31.84 0.04||48.86
REArtGS [wu2025reartgs] 0.23 13.09 46.09 25.41 55.53 28.07 0.04||63.74 0.03||45.90 0.04||54.82
@@
CD-s ArtGS [liu2025building] 85.65 3.01 2.35 1.74 10.45 20.64 1.78 0.30 1.04
REArtGS [wu2025reartgs] 45.11 2.86 14.12 2.16 15.83 16.02 0.85 0.21 0.53
@@
CD-m ArtGS [liu2025building] 216.58 9.50 0.39 7.49 0.79 46.95 0.46 11.75 6.11
REArtGS [wu2025reartgs] 71.44 10.63 119.32 70.91 136.68 81.80 0.97 10.74 40.86
@@
CD-w ArtGS [liu2025building] 1.06 4.06 2.29 3.31 9.87 4.12 1.75 0.59 1.17
REArtGS [wu2025reartgs] 3.52 2.88 13.16 6.16 17.96 8.74 0.80 0.47 0.64

### v3

@@ Abstract
Articulated objects are pervasive in daily environments, such as drawers and refrigerators. Towards their part-level surface reconstruction and joint parameter estimation, REArtGS [23] introduces a category-agnostic approach using multi-view RGB images at two different states. However, we observe that REArtGS still struggles with screw-joint or multi-part objects and lacks geometric constraints for unseen states. In this paper, we propose REArtGS++, a novel method towards generalizable articulated object reconstruction with temporal geometry constraint and planar Gaussian splatting. We first model a decoupled screw motion for each joint without type prior, and jointly optimize part-aware Gaussians with joint parameters through part motion blending. To introduce time-continuous geometric constraint for articulated modeling, we encourage Gaussians to be planar and propose a temporally consistent regularization between planar normal and depth through Taylor first-order expansion. Extensive experiments on both synthetic and real-world articulated objects demonstrate our superiority in generalizable part-level surface reconstruction and joint parameter estimation, compared to existing approaches. Project Site: https://sites.google.com/view/reartgs2/home.
@@ 1 Introduction
Articulated objects are ubiquitous in daily life, such as drawers and laptops, and their surface reconstruction holds significant value across various fields, including robotics and embodied intelligence [21, 30, 29]. Unlike static objects, articulated objects consist of multiple parts with diverse kinematic structures, requiring additional part-level mesh reconstruction and joint parameter estimation. Under this circumstance, ASDF [15] and DITTO [8] train generative models of articulated objects using expensive 3D supervisions. However, these methods exhibit limited generalization to unseen categories due to the variations in shape and motion of articulated objects. To address this limitation, PARIS [12] achieves category-agnostic articulated modeling, i.e., part-level reconstruction and motion analysis, using neural implicit radiance fields, only with multi-view RGB images from two different states. Nevertheless, PARIS faces challenges in multi-part perception and smooth dynamic surface reconstruction, suffering from the entangled implicit representation.
@@
Recently, ArtGS [13] and REArtGS [23] achieve efficient and realistic articulated object reconstruction and joint parameter estimation from two-state RGB-D and RGB images respectively, leveraging the explicit representation of 3D Gaussian Splatting (3DGS) [9]. However, we observe that these methods still suffer from the following limitations: (1) They rely on joint type priors, which require additional pipelines and pre-defined thresholds to obtain, and the joint types are assumed to be either rotation or translation. It may lead to significant errors when dealing with screw-joint and multi-part objects; (2) It is difficult to impose temporal geometric constraints from only two-state supervisions. Since the joint parameters and Gaussian primitives are jointly optimized, lacking geometric constraint for unseen states futher affects joint parameter estimation. Although REArtGS enhances geometric constraints by incorporating signed distance fields (SDF), it fails to perform temporal regularization due to the static mapping of SDF.
@@
To introduce dynamic geometric constraint for improved optimization of articulated modeling, we first compress the 3D Gaussians to flat planes to obtain accurate normals and unbiased depth maps following PGSR [1]. Then we propose a temporal geometric constraint, extending the consistent regularization of normal and depth from discrete states to entire time interval through Taylor first-order expansion. Moreover, we propose local consistent voting to address the ambiguous segmentation of overlapping regions. Experimental results of both synthetic and real-world data demonstrate our superiority over existing state-of-the-art (SOTA) methods.
@@ 2.1 Dynamic Reconstruction and Segmentation via 3DGS
3D Gaussian Spaltting [9] and its subsequent works [1, 5, 28, 14] achieve realistic and efficient scene reconstruction via explicit 3D Gaussian primitives. Recent studies [27] have increasingly explored the potential of 3D Gaussian Splatting (3DGS) for dynamic scene reconstruction. Specifically, Deformable-3DGS [9] and 4DGS [24] leverage deformation networks alongside 3D Gaussians, obtaining time-independent Gaussians for dynamic scenes. SC-GS [6] and SP-GS [2] employ superpoints for dynamic Gaussian modeling and control. Although yielding promising results, each Gaussian learns independent motion in these methods, leading to inconsistency within a rigid part when dealing with articulated objects. Moreover, they require time-continuous supervisions and face challenges in inferring motion using image collections of only two states.
Several works attempt to perform instance segmentation with 3D Gaussians [32, 16, 18, 7, 19]. These methods achieve 3D Gaussian segmentation by introducing 2D masks obtained from foundation models such as SAM [10] or DINO [31]. However, they rely on the prior knowledge provided by external models and typically struggle with part segmentation.
@@ 2.2 Articulated Object Reconstruction and Motion Analysis
Previous methods [15, 20] mainly focus on dynamic reconstruction of articulated objects. To achieve both surface reconstruction and motion analysis for articulated objects, PARIS [12, 22] employ composite neural radiance fields to jointly optimize geometry and motion through multi-view RGB images of two different states. Following this setting, several methods [3, 17, 13, 23] introduce 3DGS to articulated object reconstruction and joint parameter estimation. However, these methods require additional pipelines to obtain joint type priors, which are limited to rotation or translation, and affect the performance when dealing with screw-joint or multi-part objects.
@@ 3 Preliminary: Planar Gaussian Spaltting
3DGS [9] uses a set of Gaussians to explicitly represent the scene instead of neural implicit representation. The rendering color can be derived from the $\alpha$-blending as follows:
@@
where $\mathcal{L}_{1}$ is a pixel-level L1 norm and $\mathcal{L}_{\text{D-SSIM}}$ is the D-SSIM loss [9].
Recently, PGSR [1] obtains accurate normals and unbiased depth estimates by compressing 3D Gaussians to 2D planar representations, enhancing geometric learning. Concretely, the normal maps $\mathbf{N}$ of planar Gaussians can be naturally rendered via their shortest axes and view directions, and the unbiased depth can be estimated as:
@@ 4.1 Part-Aware Planar Gaussian Representation
Subsequently, we introduce a segmentation mask $\mathbf{M}=\left\{m_{1},m_{2},...,m_{k}\right\}$ for each planar Gaussian, representing the probability belonging to each part. Intuitively, the probability that a Gaussian primitive $\mathcal{G}_{i}$ belongs to a part is inversely proportional to its distance from the part’s center. Inspired by REACTO [20], we first denote part centers as learnable parameters $(\mathbf{O},\mathbf{V},\mathbf{\Lambda})$, where $\mathbf{O}\in\mathbb{R}^{k\times 3}$, $\mathbf{V}\in\mathbb{R}^{k\times 3\times 3}$, $\mathbf{\Lambda}\in\mathbb{R}^{k\times 3}$ are part centers, center orientations and scale vectors respectively. For Gaussian $\mathcal{G}_{i}$, the relative distances $\mathbf{L}_{i}$ between Gaussian center $\mu_{i}$ and part centers are defined as: $\mathbf{L}_{i}=\mathbf{V}(\mu_{i}-\mathbf{O})\mathbf{\Lambda}$. Then the Mahalanobis distances $\mathbf{\gamma}$ can be derived as: $\mathbf{\gamma}_{i}=\mathbf{L}_{i}^{T}\mathbf{L}_{i}$. We calculate part segmentation mask $\mathbf{M}_{i}$ of $\mathcal{G}_{i}$ through the Mahalanobis distances and a residual term learned by an multi-layer perceptron (MLP), which is illustrated in the supplementary materials.
@@ 4.2 Part-level Blending Motion
Table 1: Quantitative results on PARIS dataset. We implement all methods without depth supervision for fair comparison. Axis Pos results are measured by $m$. ”-” indicates the object containing only prismatic joints. We highlight best and second best results.
@@
Axis Ang PARIS [12] 15.72 2.02 0.03 7.43 16.62 7.17 0.71 15.40 38.28 0.03 10.34 2.90 20.16 11.53
DTA [22] 0.03 0.08 0.07 0.21 0.10 0.08 0.11 0.38 0.18 0.09 0.13 2.26 10.57 6.42
ArtGS [13] 0.01 0.03 <<0.01 0.02 <<0.01 0.01 0.02 0.04 0.02 0.01 0.02 2.03 4.56 3.30
REArtGS [23] 0.01 0.02 0.02 0.07 0.05 0.06 0.04 0.08 0.04 0.04 0.04 2.19 87.51 44.85
@@
Axis Pos PARIS [12] 10.30 9.18 0.76 5.15 15.31 35.07 13.26 24.14 - - 14.15 50.28 - 50.28
DTA [22] 1.47 1.16 2.32 1.18 1.95 1.67 1.46 4.87 - - 2.01 59.13 - 59.13
ArtGS [13] 0.06 0.43 0.94 1.35 0.18 0.38 0.15 1.06 - - 0.51 42.83 - 42.83
REArtGS [23] 6.24 0.15 1.31 3.64 0.06 0.72 1.56 0.04 - - 1.72 45.72 - 45.72
@@
Part Motion PARIS [12] 45.80 68.21 0.03 9.70 71.42 128.73 34.57 43.58 0.45 0.30 40.28 2.23 0.57 1.40
DTA [22] 0.10 0.12 0.11 0.12 0.34 0.08 0.15 0.24 <<0.01 <<0.01 0.13 2.56 0.18 1.37
ArtGS [13] 0.03 0.03 0.03 0.01 0.01 0.03 <<0.01 0.02 <<0.01 0.01 0.02 2.27 0.15 1.21
REArtGS [23] 0.31 0.23 0.06 0.08 0.06 0.15 0.05 0.04 <<0.01 0.08 0.11 38.38 0.52 19.45
@@
CD-s PARIS [12] 8.51 6.30 0.25 4.57 12.71 2.85 1.95 8.37 10.64 8.62 6.48 10.74 27.61 19.18
DTA [22] 0.27 2.48 0.64 10.06 1.64 3.08 2.64 20.73 0.46 10.73 5.74 4.50 30.98 17.74
ArtGS [13] 0.16 1.89 0.80 8.58 0.50 3.55 2.51 23.63 0.51 11.82 5.40 3.39 18.19 10.79
REArtGS [23] 0.31 1.53 0.48 6.59 0.25 1.47 0.82 16.04 0.69 7.47 3.57 7.67 22.53 15.10
@@
CD-m PARIS [12] 4.05 1.26 0.15 3.70 10.06 1.18 3.02 1.52 5.13 20.67 5.07 74.20 314.37 194.29
DTA [22] 0.46 2.27 0.96 1.36 10.25 1.61 1.37 4.45 2.07 5.34 3.01 3.63 81.36 42.50
ArtGS [13] 0.52 1.40 0.98 2.52 0.70 0.86 1.15 4.67 1.23 6.56 2.06 3.23 33.65 18.44
REArtGS [23] 0.30 0.66 0.71 1.25 0.25 0.21 0.40 0.42 1.97 12.39 1.85 2.87 175.78 89.33
@@
CD-w PARIS [12] 8.27 2.83 3.13 11.08 11.68 2.80 12.46 19.67 0.81 8.37 8.11 8.82 24.73 16.78
DTA [22] 0.58 2.36 0.75 9.35 0.65 2.95 1.43 18.68 0.51 8.34 4.56 4.47 23.48 13.98
ArtGS [13] 0.44 2.01 0.96 8.32 0.72 2.71 1.57 21.69 0.67 8.69 4.78 3.26 12.26 15.52
REArtGS [23] 0.70 1.64 0.65 10.89 0.31 3.26 1.54 17.82 0.40 7.71 4.49 2.93 11.48 7.21
@@ 4.5 Initialization and Optimization
Inspired by ArtGS, we first perform a fast optimization for two sets of Gaussians at both start and end states using vanilla 3DGS pipeline [9], then identify the dynamic Gaussians for each state by criterion: $\Delta\textbf{x}_{i}>\tau\textbf{x}$, where $\Delta\textbf{x}_{i}$ is the Chamfer Distance (CD) from Gaussian $\mathcal{G}_{i}$ at one state to all Gaussians of another state. x is the maximum CD and $\tau$ is set to 0.02. We initialize static and dynamic Gaussians of canonical state as the mean of the two states, using Hungarian matching.
@@ 4.6 Dynamic Mesh Extraction
Table 2: Quantitative results on ArtGS-Multi dataset with additional 2 objects exhibiting screw joints. Due to the large number of parts, we report the average metric for all movable parts. We implement all methods without depth supervision for fair comparison. Part motion error for screw objects is decomposed by translation||rotation. Axis Pos results are measured by $m$. ‘-’ indicates the object contains only prismatic joints. We highlight best and second best results.
@@
Axis Ang ArtGS [13] 1.23 83.69 0.01 34.83 5.62 25.08 42.28 87.57 64.93
REArtGS [23] 85.45 42.18 64.34 30.69 24.50 49.43 46.45 87.80 67.13
@@
Axis Pos ArtGS [13] - 70.93 0.36 235.19 8.54 96.49 50.64 38.75 44.70
REArtGS [23] - 155.80 500.51 31.44 233.48 230.31 47.18 20.41 33.80
@@
Part Motion ArtGS [13] 0.08 22.50 0.01 18.75 5.73 9.41 0.06||65.87 0.02||31.84 0.04||48.86
REArtGS [23] 0.23 13.09 46.09 25.41 55.53 28.07 0.04||63.74 0.03||45.90 0.04||54.82
@@
CD-s ArtGS [13] 85.65 3.01 2.35 1.74 10.45 20.64 1.78 0.30 1.04
REArtGS [23] 45.11 2.86 14.12 2.16 15.83 16.02 0.85 0.21 0.53
@@
CD-m ArtGS [13] 216.58 9.50 0.39 7.49 0.79 46.95 0.46 11.75 6.11
REArtGS [23] 71.44 10.63 119.32 70.91 136.68 81.80 0.97 10.74 40.86
@@
CD-w ArtGS [13] 1.06 4.06 2.29 3.31 9.87 4.12 1.75 0.59 1.17
REArtGS [23] 3.52 2.88 13.16 6.16 17.96 8.74 0.80 0.47 0.64